# Unbiased Single-Queried Gradient for Combinatorial Objective

### v1

@@ 2.2 Continuous relaxations and straight-through gradients
Recent papers (Geng et al., 2025; Li et al., 2025) have suggested that a relaxation to smooth value may work as long as $\nabla V(x)$ exists pathwise almost everywhere. However, the paper (Sornwanee, 2026b) shows that it is incorrect. Even in 1 dimensional case, the interchange for
@@
This is generally not true when $f$ is discontinuous, which can happen when thresholding333which is similar to straight-through estimator where we have to select one discrete variable, often the hard maximum. This problem is then similar to the local optimization method in multi-level optimizations (Jalota et al., 2024; Sornwanee, 2026a). is done. In our design, we will use a function such that, at the thresholding boundary, the value is always 0 to avoid the discontinuity implication.
@@ 2.3 Zeroth-order / black-box gradient methods
When only function values are available, classical zeroth-order methods approximate gradients via finite differences, dating back to Kiefer-Wolfowitz stochastic approximation (Kiefer and Wolfowitz, 1952). Simultaneous Perturbation Stochastic Approximation (Spall, 1992) reduces query complexity by perturbing all coordinates simultaneously, and related random-direction estimators form the basis of modern derivative-free optimization. These methods typically require smoothness assumptions on the underlying objective and may still require multiple oracle calls per update for competitive variance. (Bengio et al., 2013) In contrast, our focus is on expectations induced by discrete random mappings $K(x)$ and on estimators tailored to this structure, with the goal of minimizing oracle calls.
@@ 3 Stochastic Gradient
Note that, we require almost surely $\omega\in\Omega$ after fixing a value of $x\in(0,1)^{d}$ since, conditioned444in a do-algebra sense on the arbitrary current value of $x\in(0,1)^{d}$, we want to be able to retrieve the stochastic gradient with probability $1$.
@@ 4.2 Variance, Calibration, & Importance Sampling
For example, in a one-dimensional case, if $Q(0)=0$ and $Q(1)=1$, we will have that the ground truth gradient $g(x)=1$. Both REINFORCE gradient and ESG: LongJump are unbiased. However, the variance of REINFORCE gradient is $\frac{1-x}{x}$, which explodes to $\infty$ when $x\to 0$, while the gradient obtained via the ESG algorithm is 11.888The use of pathwise gradient may be responsible for the variance reduction in some regime. (Asmussen and Glynn, 2007)
@@ 5.2 Encoded ESG & Encoded SQD
These encoded versions [^12] See the appendix A for the encoded ESG algorithm 3 and the ecnoded SQD algorithm 4. may reduce the autodifferentiation work load since the variable $z$ is linear in $e$. Moreover, the variance may be reduced when $x$ is close to the boundary since we do not require the division by $\hat{\sigma}^{\prime}\left(\hat{\sigma}^{-1}(x_{i})\right)$. From the experiment131313as shown in the appendix E, we observe that the trajectory tends to be more gradual.
@@ 6 Conclusion
Acknowledgement
The author thanks Rajat Dwaraknath, Haozhan Gao, Zijie Geng, Xuchen Gong, Lutong Hao, Senem Isik, Jiwon Jang, Chartsiri Jirachotkulthorn, Minseo Kim, Anish Senapati, Ellen Vitercik, Jenny Xu, and Yunzhi Zhang for useful discussions.

### v2

@@ 2.2 Continuous relaxations and straight-through gradients
Recent papers (Geng et al., 2025; Li et al., 2025) have suggested that a relaxation to smooth value may work as long as $\nabla V(x)$ exists pathwise almost everywhere. However, the paper (Sornwanee, 2026) shows that it is incorrect. Even in 1 dimensional case, the interchange for
@@
This is generally not true when $f$ is discontinuous, which can happen when thresholding333which is similar to straight-through estimator where we have to select one discrete variable, often the hard maximum. This problem is then similar to the local optimization method in multi-level optimizations (Jalota et al., 2024; Sornwanee, 2025). is done. In our design, we will use a function such that, at the thresholding boundary, the value is always 0 to avoid the discontinuity implication.
@@ 2.3 Zeroth-order / black-box gradient methods
When only function values are available, classical zeroth-order methods approximate gradients via finite differences, dating back to Kiefer-Wolfowitz stochastic approximation (Kiefer and Wolfowitz, 1952). Simultaneous Perturbation Stochastic Approximation (Spall, 1992) reduces query complexity by perturbing all coordinates simultaneously, and related random-direction estimators form the basis of modern derivative-free optimization. These methods typically require some smoothness assumptions [^4] Even in a discrete grid case, the papers (Wang and Spall, 2011; Wang et al., 2025) require the approximation of the gradient by the midpoint gradient. on the underlying objective and may still require multiple oracle calls per update for competitive variance. (Bengio et al., 2013) In contrast, our focus is on expectations induced by discrete random mappings $K(x)$ and on estimators tailored to this structure, with the goal of minimizing oracle calls.
@@ 3 Stochastic Gradient
Note that, we require almost surely $\omega\in\Omega$ after fixing a value of $x\in(0,1)^{d}$ since, conditioned555in a do-algebra sense on the arbitrary current value of $x\in(0,1)^{d}$, we want to be able to retrieve the stochastic gradient with probability $1$.
@@ 4.2 Variance, Calibration, & Importance Sampling
For example, in a one-dimensional case, if $Q(0)=0$ and $Q(1)=1$, we will have that the ground truth gradient $g(x)=1$. Both REINFORCE gradient and ESG: LongJump are unbiased. However, the variance of REINFORCE gradient is $\frac{1-x}{x}$, which explodes to $\infty$ when $x\to 0$, while the gradient obtained via the ESG algorithm is 11.999The use of pathwise gradient may be responsible for the variance reduction in some regime. (Asmussen and Glynn, 2007)
@@ 5.2 Encoded ESG & Encoded SQD
These encoded versions [^13] See the appendix A for the encoded ESG algorithm 3 and the ecnoded SQD algorithm 4. may reduce the autodifferentiation work load since the variable $z$ is linear in $e$. Moreover, the variance may be reduced when $x$ is close to the boundary since we do not require the division by $\hat{\sigma}^{\prime}\left(\hat{\sigma}^{-1}(x_{i})\right)$. From the experiment141414as shown in the appendix E, we observe that the trajectory tends to be more gradual.
@@ 6 Conclusion
Acknowledgements
The author thanks Rajat Dwaraknath, Haozhan Gao, Zijie Geng, Xuchen Gong, Lutong Hao, Senem Isik, Jiwon Jang, Chartsiri Jirachotkulthorn, Minseo Kim, Anish Senapati, James Spall, Ellen Vitercik, Jenny Xu, and Yunzhi Zhang for useful discussions.